# Table-R1: Inference-Time Scaling for Table Reasoning

### v1

@@ Abstract
Zheyuan Yangthanks: Equal Contributions.  Lyuhao Chen  Arman Cohan  Yilun Zhao Yale NLP Lab
@@ 2.2 Table Reasoning
Reasoning over tabular data has long attracted attention due to its practical applications in real-world scenarios such as data analysis. It encompasses a variety of tasks, including short-form question answering Pasupat and Liang (2015); Cheng et al. (2022); Lu et al. (2023); Zhao et al. (2023d, 2024b); Wu et al. (2025), fact verification Chen et al. (2020); Gupta et al. (2020), and free-form question answering Nan et al. (2022); Zhao et al. (2023a). Earlier research primarily focused on fine-tuning smaller language models for specific tasks Herzig et al. (2020); Zhao et al. (2022); Liu et al. (2022). More recently, efforts such as TableLlama Zhang et al. (2024a), TableLLM Zhang et al. (2025a), and the TableGPT series Zha et al. (2023); Su et al. (2024) have advanced the adaptation of LLMs for table reasoning, enabling more general-purpose capabilities across task types. In parallel, several studies have begun exploring agentic approaches to tackle table reasoning tasks Ye et al. (2023); Zhao et al. (2024a); Nan et al. (2024); Yu et al. (2025a). Despite these advances, the application and enhancement of inference-time scaling for table reasoning remain largely underexplored. Our study shows that 7B-scale LLMs with inference-time scaling can match the performance of frontier models such as GPT-4.1. This finding highlights a promising direction for advancing table reasoning.
@@ 3.1 Training Data Collection
We construct the Table-R1 training dataset by integrating three representative table reasoning tasks, each introducing distinct reasoning challenges: (1) Short-form Table QA, which requires models to provide precise answers to questions grounded in tabular data; (2) Table Fact Verification, which requires models to determine whether a given claim is entailed by the table content; and (3) Free-form Table QA, which requires models to produce open-ended answers grounded in tabular information. Each task contributes unique reasoning challenges, and we select established datasets (presented in Table 1) to ensure comprehensive coverage. All datasets are preprocessed with the presence of verifiable ground truths for reward computation to align with the RLVR paradigm.
@@ 3.2 Table-R1-SFT via Supervised Finetuning
To enable inference-time scaling in Table-R1-SFT, we curate a new table reasoning dataset featuring long CoT reasoning. Specifically, for each instance in the raw training data described in the previous subsection, we use DeepSeek-R1 DeepSeek-AI et al. (2025) to generate a long CoT response. The response consists of a step-by-step reasoning process followed by a final answer. We present the prompts for response generation in Appendix C.3. To ensure the quality and correctness of the training data, we apply automated evaluators (detailed in Section 4.2) to assess the final answers. Examples with incorrect answers are filtered out. After this verification step, we obtain a total of 33,601 high-quality examples for SFT training. This dataset is then used to train the Table-R1-SFT model.
@@ 4.2 Experiment Setup
To address the lack of a unified evaluation framework for table reasoning, we introduce a new benchmark suite encompassing a wide range of datasets. For in-domain evaluation, we use test sets from the same distributions as the training data: WTQ Pasupat and Liang (2015) and HiTab Cheng et al. (2022) for TQA; TabFact Chen et al. (2020) for TFV; and FeTaQA Nan et al. (2022) for FF-TQA. To assess generalization of Table-R1, we further conduct out-of-domain evaluation on datasets that are not seen during training: TabMCQ Jauhar et al. (2016), TabMWP Lu et al. (2023), and FinQA Chen et al. (2021) for TQA; InfoTabs Gupta et al. (2020), PubHealthTab Akhtar et al. (2022), and Feverous Aly et al. (2021) for TFV; ToTTo Parikh et al. (2020), QTSumm Zhao et al. (2023b), and RotoWire Wiseman et al. (2017) for FF-TQA. We provide detailed descriptions of the evaluated baseline systems in Appendix B.
@@
For each evaluated dataset, we use its test set for evaluation. For the TQA task, we report Exact Match (EM) accuracy. For TFV, we use classification accuracy. For FF-TQA, we measure generation quality with BLEU and ROUGE-L scores. Considering that EM accuracy in short-answer TQA may underestimate model performance due to formatting variations or semantically equivalent but non-exact matches, we further re-evaluate responses initially marked incorrect by EM using the GPT-4.1 mini model, with the prompt shown in Appendix C.3.
[…]
@@ 4.3 Main Findings
[…]
@@ 4.5 Reasoning Capacity Boundary
Model In-domain Performance Out-of-domain Performance
FF-TQA TFV TQA FF-TQA TFV TQA
Model-level Ablation
Base vs. Instruct
[…]
Task-level Ablation
Cross-task Generalization
Format Ablation
Effect of Format Reward
@@ 4.6 Qualitative Analysis
To gain deeper insights into how RLVR shapes model behavior, we conduct a qualitative analysis of model responses throughout the training process. By examining the same set of representative prompts across different RL training steps, we observe that the model not only internalizes general R1-style reasoning characteristic, but also acquires table-specific reasoning abilities critical for tabular tasks, as illustrated in Figure 6 and Appendix D.
@@ 4.7 Ablation Studies on RLVR Training
[…]
@@
We assess the role of explicit reasoning format supervision by removing the format reward during training. This ablation reduces training stability and slightly lowers in-domain performance. More notably, generalization suffers: while TFV scores may improve, performance on short-answer TQA and FF-TQA declines, sometimes even below baseline, indicating that format supervision is crucial for transferable reasoning. We also test inference-time control by restricting the initial output token to <think> or <answer>. However, this constraint does not alter the model’s internal reasoning or task performance, showing that inference-time decoding tricks cannot substitute for training with flexible reasoning formats.
@@ Acknowledgments
Appendix A Information of Evaluated Table Reasoning Datasets
@@ B.1 Baseline Systems
We benchmark our approach against a comprehensive set of strong baselines, encompassing both proprietary and open-source models. Among proprietary models, we include GPT-4.1 and GPT-4.1 mini. For open-source baselines, we evaluate general purpose LLMs from the Qwen2.5 and LLaMA3 series, reasoning models such as Deepseek-R1, its official distilled variant, our SFT model distilled from Deepseek-R1 on table reasoning data (as described in Section 3.2), and QwQ-32B, as well as table-oriented models including TableLlama, TableLLM, and TableBenchLLM. Additional evaluation details, including specific model configurations and prompt templates, are provided in Appendix C.
Appendix C Evaluation Details
Additional evaluation details, including inference setup, evaluated model configurations and prompt templates.
C.1 Inference Setup
@@
C.2 Evaluated Model Configuration
@@ C.3 Prompts
    "answer": [
        "answer1",
        "answer2",
@@
Appendix D Qualitative Analysis Cases

### v2

@@ Abstract
Zheyuan Yangthanks: Equal Contributions.1   Lyuhao Chen∗2   Arman Cohan1   Yilun Zhao1 1Yale NLP Lab  2Carnegie Mellon University
@@ 2.2 Table Reasoning
Reasoning over tabular data has long attracted attention due to its practical applications in real-world scenarios such as data analysis. It encompasses a variety of tasks, including short-form question answering Pasupat and Liang (2015); Cheng et al. (2022); Lu et al. (2023); Zhao et al. (2023d, 2024b); Wu et al. (2025), fact verification Chen et al. (2020); Gupta et al. (2020), and free-form question answering Nan et al. (2022); Zhao et al. (2023a). Earlier research primarily focused on fine-tuning smaller language models for specific tasks Herzig et al. (2020); Zhao et al. (2022); Liu et al. (2022). Recent efforts have advanced the adaptation of LLMs for table reasoning Zhang et al. (2024a, 2025a); Su et al. (2024); Zha et al. (2023); Su et al. (2024); Deng et al. (2025), enabling more general-purpose capabilities across task types. In parallel, several studies have begun exploring agentic approaches to tackle table reasoning tasks Ye et al. (2023); Zhao et al. (2024a); Nan et al. (2024); Yu et al. (2025a). Despite these advances, the application and enhancement of inference-time scaling for table reasoning remain largely underexplored. Our study shows that 7B-scale LLMs with inference-time scaling can match the performance of frontier models such as GPT-4.1. This finding highlights a promising direction for advancing table reasoning.
@@ 3.1 Training Data Collection
We construct the Table-R1 training dataset by integrating three representative table reasoning tasks, each introducing distinct reasoning challenges: (1) Short-form Table QA, which requires models to provide precise answers to questions grounded in tabular data; (2) Table Fact Verification, which requires models to determine whether a given claim is entailed by the table content; and (3) Free-form Table QA, which requires models to produce open-ended answers grounded in tabular information. Each task contributes unique reasoning challenges, and we select established datasets (presented in Table 1) to ensure comprehensive coverage. All datasets are preprocessed with the presence of verifiable ground truths for reward computation to align with the RLVR paradigm. More detailed explanation of training data is presented in Appendix A.1.
@@ 3.2 Table-R1-SFT via Supervised Finetuning
To enable inference-time scaling in Table-R1-SFT, we curate a new table reasoning dataset featuring long CoT reasoning. Specifically, for each instance in the raw training data described in the previous subsection, we use DeepSeek-R1 DeepSeek-AI et al. (2025) to generate a long CoT response. The response consists of a step-by-step reasoning process followed by a final answer. We present the prompts for response generation in Appendix C. To ensure the quality and correctness of the training data, we apply automated evaluators (detailed in Section 4.2) to assess the final answers. Examples with incorrect answers are filtered out. After this verification step, we obtain a total of 33,601 high-quality examples for SFT training. This dataset is then used to train the Table-R1-SFT model.
@@ 4.2 Experiment Setup
To address the lack of a unified evaluation framework for table reasoning, we introduce a new benchmark suite encompassing a wide range of datasets. For in-domain evaluation, we use test sets from the same distributions as the training data: WTQ Pasupat and Liang (2015) and HiTab Cheng et al. (2022) for TQA; TabFact Chen et al. (2020) for TFV; and FeTaQA Nan et al. (2022) for FF-TQA. To assess generalization of Table-R1, we further conduct out-of-domain evaluation on datasets that are not seen during training: TabMCQ Jauhar et al. (2016), TabMWP Lu et al. (2023), and FinQA Chen et al. (2021) for TQA; InfoTabs Gupta et al. (2020), PubHealthTab Akhtar et al. (2022), and Feverous Aly et al. (2021) for TFV; ToTTo Parikh et al. (2020), QTSumm Zhao et al. (2023b), and RotoWire Wiseman et al. (2017) for FF-TQA. We provide detailed descriptions of the evaluated datasets and evaluated baseline systems in Appendix A.2 and Appendix B.
@@
For each evaluated dataset, we use its test set for evaluation. For the TQA task, we report Exact Match (EM) accuracy. For TFV, we use classification accuracy. For FF-TQA, we measure generation quality with BLEU and ROUGE-L scores. Considering that EM accuracy in short-answer TQA may underestimate model performance due to formatting variations or semantically equivalent but non-exact matches, we further re-evaluate responses initially marked incorrect by EM using the GPT-4.1 mini model, with the prompt shown in Appendix C.
ȷ RQ1: How effective is inference-time scaling for table reasoning tasks?
@@ 4.3 Main Findings
ȷ RQ2: What contributes to the success of RLVR methods in table reasoning tasks?
@@ 4.6 Qualitative Analysis
To gain deeper insights into how RLVR shapes model behavior, we conduct a qualitative analysis of model responses throughout the training process. By examining the same set of representative prompts across different RL training steps, we observe that the model not only internalizes general R1-style reasoning characteristic, but also acquires table-specific reasoning abilities critical for tabular tasks, as illustrated in Figure 5 and Appendix D.1.
@@ 4.7 Ablation Studies on RLVR Training
ȷ RQ3: How do various factors influence the effectiveness of RLVR training in table reasoning tasks?
@@
We assess the role of explicit reasoning format supervision by removing the format reward during training. This ablation reduces training stability and slightly lowers in-domain performance. More notably, generalization suffers: while TFV scores may improve, performance on short-answer TQA and FF-TQA declines, sometimes even below baseline, indicating that format supervision is crucial for transferable reasoning.
@@ Acknowledgments
Appendix A Dataset Details
A.1 Detailed Explanation of Training Data
Our training data originates from four benchmark datasets: WTQ, HiTab, TabFact, and FeTaQA.
Training Data for RLVR. For WTQ, HiTab, and FeTaQA, we used their complete training sets. For TabFact, which has a very large training set (>90,000>90,000 samples), we implemented a difficulty-based filtering process to create a more focused and challenging subset. Specifically, we used Qwen2.5-7B-Instruct to generate $16$ candidate responses for each sample. We retained all samples with a pass rate of $\leq 7/16$ and randomly sampled $10,000$ instances from those with a pass rate between $8/16$ and $15/16$. This resulted in a final set of $20,740$ samples for TabFact. After constructing the prompts, we filtered out any instances where the prompt length exceeded $4096$ tokens. This final, combined dataset ($48,563$ samples) was used for RLVR training.
Training Data for SFT. We used the RLVR dataset as a starting point. We then generated reasoning traces for each sample using DeepSeek-R1 and applied a stringent quality-control filter to retain only high-quality examples. For WTQ, HiTab, and TabFact, we kept only the samples where the generated answer was an exact match to the ground truth. For FeTaQA, we kept samples where the average of BLEU and ROUGE-L scores was $\geq 0.35$. This resulted in a high-quality SFT dataset of $33,601$ samples.
The data volumes at each stage are summarized below, as shown in Table 3.
Dataset Raw Samples Samples for RLVR Samples for SFT (Filtered)
WTQ 14,152 13,706 9,524
HiTab 7,301 6,793 4,170
TabFact 92,283 20,740 16,006
FeTaQA 7,326 7,324 3,901
Total 121,062 48,563 33,601
Table 3: The data volumes at each stage.
Table 4 provides detailed statistics on the SFT data. Notably, the quality filtering significantly reduces the average response-to-prompt length ratio (from $1.73$ to $1.49$ overall). This aligns with the common observation that incorrect or lower-quality reasoning traces are often longer and more convoluted, reinforcing the effectiveness of our filtering strategy.
Dataset / Length (Chars) Min Prompt Max Prompt Avg. Prompt Min Resp. Max Resp. Avg. Resp. Avg. Resp/Prompt Ratio (Before/After Filtering)
WTQ 625 14,192 2,114.67 465 28,023 2,356.26 1.78 / 1.45
HiTab 694 10,264 3,251.17 526 32,630 2,827.42 1.57 / 1.12
TabFact 624 8,906 1,767.64 529 28,762 2,474.04 1.76 / 1.61
FeTaQA 473 13,793 1,656.14 546 28,661 1,982.57 1.71 / 1.53
Overall 473 14,192 2,037.17 465 32,630 2,427.45 1.73 / 1.49
Table 4: Detailed statistics on the SFT data.
A.2 Information of Evaluated Table Reasoning Datasets
@@ B.1 Baseline Systems
We benchmark our approach against a comprehensive set of strong baselines, encompassing both proprietary and open-source models. Among proprietary models, we include GPT-4.1 and GPT-4.1 mini. For open-source baselines, we evaluate general purpose LLMs from the Qwen2.5 and LLaMA3 series, reasoning models such as Deepseek-R1, its official distilled variant, our SFT model distilled from Deepseek-R1 on table reasoning data (as described in Section 3.2), and QwQ-32B, as well as table-oriented models including TableLlama, TableLLM, and TableBenchLLM.
B.2 Inference Setup
@@
B.3 Evaluated Model Configuration
@@ Appendix C Prompts
    "answer": ["answer1", "answer2", ...]
@@
Appendix D Analysis
D.1 Qualitative Analysis Cases
@@
D.2 Reasoning Capacity Boundary
Figure 8: Pass@k performance on WTQ and HiTab.
D.3 Ablation Studies
Model In-domain Performance Out-of-domain Performance
$F$-TQA TFV TQA $F$-TQA TFV TQA
Model-level Ablation
Base vs. Instruct
[.4pt/2pt]   Model Architecture Comparison
[.4pt/2pt]   SFT on Domain-Specific Data
Task-level Ablation
Cross-task Generalization
Format Ablation
Effect of Format Reward